# Constraining the Stellar-to-Halo Mass Relation with Galaxy Clustering and Weak Lensing from DES Year 3 Data

### v2

@@ 4.1.2 The stellar mass sample
With the GalPro output from our approximately complete galaxy sample, we now define the stellar mass sample for this work. In particular, we like to define bins in both redshift and stellar mass, as we are interested in the redshift-dependence of the SHMR. For the redshift binning, we define the binning similar to that used in DES Y3 (e.g. Prat et al., 2022; Abbott et al., 2022). That is, we bin the galaxies into four tomographic redshift bins based on their mean redshift $z_{\rm m}$, as shown in Table 1. For each redshift bin, we first impose a stellar mass cut on the mean stellar mass $M_{\star,\rm m}$ so that the remaining galaxies are complete in stellar mass at that redshift [^1] This is done using a similar procedure as that is described in Section 5.2 in Pozzetti et al. (2010). For the four redshift bins we get $\log M_{\rm lim}=(9.5617,9.9996,10.3549,10.5923)$.. We then split the remaining galaxies into stellar mass bins again according to their mean stellar mass $M_{\star,\rm m}$ where aside from the highest mass bins, each bin contains $3\times 10^{6}$ objects. The number of objects, number density, as well as the limits in stellar-mass per bin are shown in Table 1. From now on, we will denote a single bin with the indices $(\ell,m)$, the first of which refers to the reshift bin and the latter to the stellar-mass bins in that $\ell$. The approximate redshift bin edges are defined, per $\ell$ index, as $\{(0.20,0.40),(0.40,0.55),(0.55,0.70),(0.70,0.85)\}$ for =$\ell=\{1,2,3,4\}$ respectively.
@@
($\log M_{\star}/M_{\odot}$) [×$[\times 10^{6}]$ $[{\rm arcmin^{-2}}]$ $[M_{\odot}]$
@@ 4.1.3 Weights
Thus far, we have constructed a sample of galaxies that is complete in stellar mass and binned by stellar mass and redshift. However, in most cases, the rather simplistic selection will still leave low-level spatial patterns in the galaxy distributions that are imprints of observing conditions such as depth or seeing. A common way to further remove these effects is to assign each galaxy a weight to compensate the over- or under-estimation of the galaxy densities – we will refer these weights as large-scale structure (LSS) weights (see also Rodríguez-Monroy et al., 2022; Sánchez et al., 2022).
@@
The methodology we implement in this work is very similar to what is done in Sánchez et al. (2022). It relies on using a neural network to find the mapping between the galaxy number density and a number of survey conditions, and then correcting for them, with the underlying assumption that the cosmological LSS signal should not be correlated with observing conditions. In particular, we examine the following observational parameters via a set of survey property (SP) maps, each constructed in all four $griz$ bands:
@@
  • Galactic extinction: We use the SFD dust extinction map from Schlegel et al. (1998), which measures the $E(B-V)$ reddening, in magnitudes.
  • Stellar density: We use a map of stellar density, in ${\rm deg}^{-2}$, using stellar sources from Gaia EDR3 (Gaia Collaboration, 2021).
@@ 4.2 Source galaxies
We use the DES Y3 shear catalog presented in Gatti, Sheldon et al. (2020). The galaxy shapes are estimated using the Metacalibration (Huff & Mandelbaum, 2017; Sheldon & Huff, 2017) algorithm. The shear catalog has been thoroughly tested in Gatti, Sheldon et al. (2020), and tests specifically tailored for tangential shear have been presented in Prat et al. (2022). In this paper, we perform additional tests on this shear catalog for tangential shear measurement on small scales (Section 5.3).
We further note that, it was recently discovered that the Year 3 DES cosmological inference results in Abbott et al. (2022) used a tomographic binning of the source catalog that was inconsistent with the redshift distribution of the galaxies. The first iteration of our results also suffered from this inconsistency but we have since updated all measurements and our Bayesian analysis to use the correct version of the source catalog, also used in McCullough et al. (2024). Comparing the two analyses, we found mostly negligible effects from updating the catalog, although some of the trends in our host halo mass, galaxy bias and satellite fraction constraints changed as a function of stellar mass and redshift. We only present results using the correct version of the catalog in this current paper.
Following Prat et al. (2022) we bin the source galaxies into four redshift bins, where details of the redshift binning and calibration is described in Myles, Alarcon et al. (2020). The redshift distributions of the sources is shown in Figure 3. The number of galaxies per source bin and the galaxy number density per square arcmin are both listed in Table 1. The approximate redshift bin edges, per $s$ index, defined as $\{(0.20,0.43),(0.43,0.63),(0.63,0.90),(0.90,1.30)\}$ for $s=\{1,2,3,4\}$ respectively.
@@ 5.1 Data vectors
The galaxy-galaxy lensing estimator is the same as in Prat et al. (2022) and written as
@@
For galaxy clustering we adopt the Landy-Szalay (Landy & Szalay, 1993) estimator, written as:
@@
For both $\gamma_{t}(\theta)$ and $w(\theta)$ we use $N_{\theta}=30$ angular bins in the range of $[0.25,250]\;{\rm arcmin}$. Each galaxy $i$ in our sample is weighted by $w_{g,i}$ as described in Section 4.1.3. Our catalog of random points contains $30$ times the number of galaxies in our lens galaxy catalog.
@@ 5.2 Covariance matrix
where $\gamma_{t}^{k}(\theta_{i})$ is the shear in the $i$’th angular bin for the $k$’th JK resampling, $\langle\gamma_{t}(\theta_{i})\rangle_{k}$ is the average over all $N_{\rm JK}$ realizations of the shear for the $i$’th angular bin and we have defined $\Delta\gamma^{k}_{i}\equiv\gamma_{t}^{k}(\theta_{i})-\langle\gamma_{t}(\theta_{i})\rangle_{k}$. We use $N_{\rm JK}=200$ JK patches defined via the kmeans [^2] https://github.com/esheldon/kmeans__\__radec algorithm. $N_{\rm JK}$ is chosen so that the individual JK regions are at least as large as the maximum angular scale we need for our measurements.
When inverting the covariance matrix in the likelihood analysis, a correction factor is needed to account for the bias introduced from the noisy covariance (Friedrich et al., 2016). This correction is often referred to as the Hartlap (Hartlap et al., 2007) correction. When inverting the JK covariance matrix $\mathcal{C}$ we multiply it by a factor $H$ to get the unbiased covariance (Kaufman, 1967)
@@
where $N_{\theta}$ is the number of angular bins we use in each model fit. As shown in Hartlap et al. (2007), for $N_{\theta}/N_{\rm JK}<0.8$ the correction produces an unbiased estimate of the inverse covariance matrix; in our case we find that, depending on the bin we consider, $N_{\theta}/N_{\rm JK}$ can vary between $0.215$ and $0.240$. In Appendix D we also compare our Jackknife covariance with an analytical Gaussian covariance calculated via the software package CosmoCov (Krause & Eifler, 2017; Fang et al., 2020) to show that they are fairly consistent and that our main results should not be sensitive to our choice of how we model the covariance matrix.
@@ 5.4 Bin selection and scale cuts
$\Delta z_{s}^{3}$ $\mathcal{N}(0,0.01)$ Source photo-$z$ uncertainty in $3$rd bin; eq. (32)
@@
$m_{s}^{3}$ $\mathcal{N}(-0.0255,0.0013)$ Multiplicative bias in $3$rd source bin; eq. (27)
@@ 6 Parameter inference
We continue to perform parameter inference by fitting our model described in Sections 2 and 3 to the data vector described in Section 5. In our fiducial model, we vary the seven SHMR parameter vector $\boldsymbol{\theta}_{\rm SHMR}=(\log M_{1},\log\epsilon,\alpha,\gamma,\delta,B_{\rm sat},B_{\rm cut},\beta_{\rm sat},\beta_{\rm cut})$. We also vary three HOD parameters $\boldsymbol{\theta}_{\rm HOD}=(\sigma_{\log M_{\star}},\alpha_{\rm sat},\kappa)$, as well as the additional model-complexity parameters $\boldsymbol{\theta}_{\rm complex}=(\kappa_{c},\alpha_{\rm lmag})$. Since the distribution of satellite galaxies is typically less concentrated than that of the dark matter (e.g. Carlberg et al., 1997; Lin et al., 2004; Nagai & Kravtsov, 2005; Hansen et al., 2005), we vary $\kappa_{c}$ within a range mostly below unity. For the magnification parameter $\alpha_{\rm lamg}$ we use values between $1$ (for no magnification) and $3$ in order to be broadly consistent with the constraints for the MagLim sample from DES Y3 (Elvin-Poole et al., 2023). Given that our samples are expected to be complete in our redshift-stellar mass bins, we decided to set the central and satellite parameters $f_{\rm cen}$ and $f_{\rm sat}$, respectively, to unity.
@@
For observational systematic effects, we further vary a set of systematic parameters for each lens-source combination $\boldsymbol{\theta}_{\rm sys}=(m_{s}^{j},\Sigma_{\ell}^{i},\Delta z_{\ell}^{i},\Delta z_{s}^{j},\Delta\log M_{\star}^{i})$ for lens bin $i$ and source bin $j$. The functions for all these uncertainty parameters are defined in Section 3.5. To the source-bin systematics parameters $\Delta z_{s}$ and $m_{s}$ we apply the same Gaussian priors as in the DES Y3 main analyses (see Myles et al., 2021; MacCrann et al., 2022). The priors on the lens-bin redshift uncertainty parameters $\Delta z_{\ell}$ and $\Sigma_{\ell}$ are set so that they are broadly consistent with the MagLim sample from DES Y3. More specifically, we apply a Gaussian prior on $\Delta z_{\ell}$ centered at 00 with a variance of $0.1$ and a uniform prior from $0.8$ to $1.3$ to the stretch parameter $\Sigma_{\ell}$. Finally, we allow the stellar-mass distribution of our lenses to have a small shift, $\Delta\log M_{\star}$, to which we apply uniform priors from $-0.15$ to $0.15$, to account for about $0.3$ dex uncertainty on the mean stellar mass which is typical for inference of this parameter from SED fitting algorithms. This is a fairly conservative choice given our validation tests on the stellar mass catalog (Appendix B).
@@
The priors on the lens magnification parameter $\alpha_{\rm lmag}$ is allowed to vary from $1$, which translates to no magnification, to $3$, which gives a lot of freedom to this parameter component. The parameter $\kappa_{c}$, described in Section 3.6, is allowed to mostly get values less than unity since the satellite galaxies are expected to be less concentrated than central. The SHMR parameter priors are informed by Kravtsov et al. (2018) and allow for considerable freedom during our model fits. As for the HOD parameters, we set broad priors similar to relevant studies, such as Leauthaud et al. (2011); Zu & Mandelbaum (2015).
@@
To sample the posterior of each data set we utilize the Multinest [^3] https://github.com/JohannesBuchner/MultiNest sampler, which implements a nested sampling algorithm (see for example Feroz et al., 2009), and assume a Gaussian likelihood. For data vector $\mathbf{d}$ of length $N_{d}$ and model prediction vector $\mathbf{m}$ of the same length we express the log-likelihood as
@@
where $\mathcal{C}_{\Pi}$ and $\mathcal{C}_{p}$ are the prior and posterior covariance, respectively. We should note here that a goodness-of-fit estimation based on finding an effective number of parameters is not always straightforward when the parameters do not enter the model linearly, as discussed in Section 6.3 of Joachimi et al. (2021).
@@ 7 Results
We now present the main results from this work – the model fits to our data as well as constraints on the physical quantities of our model (galaxy bias, satellite fraction, and the SHMR). We first discuss our fiducial results in Section 7.1 and then explore some of the dependencies of our model fit to some analysis choices in Section 7.2. We use our pipeline to jointly fit lensing and galaxy clustering. As we have previously shown in Zacharegkas et al. (2022), our HOD model is capable of describing galaxy-galaxy lensing well. The two main advances compared to that work are: 1) we are using the stellar mass-selected generic galaxy sample described in Section 5, and 2) we include galaxy clustering as part of the model fit.
@@ 7.1 Constraints from DES Y3 galaxy clustering and galaxy-galaxy lensing
For all the fits, we show the model broken down into the 1- and 2-halo, central and satellite terms. We find that our model generally provides a good enough description of the data points, in all lens and source pairs and on all scales, to capture the main features of the dataset. However, our model is not a perfect fit to the data, especially for $w(\theta)$. In particular, we find that to fit the 1-halo to 2-halo transition regime for low-mass low-redshift bins, it is necessary to use the prescription described in Section 3.3 instead of the model used in Zacharegkas et al. (2022).
We have used the reduced $\chi^{2}$ values, which we list in Table 1, from fitting our model to the data as a goodness-of-fit guide when determining how well the fits perform using different iterations of our fiducial model. The goodness-of-fit numbers we obtain are much larger than unity per degree of freedom, which, at first glance, implies a poor fit of the model to the data. In principle, we could add additional complexities (see Section 3.6) to the model to make the $\chi^{2}/N_{\rm dof}$ closer to 1, but that would create additional degeneracies that can be difficult to interpret. Rather, we keep the most relevant and physically motivated model components here to capture the qualitative behaviors of the data. In future work, the basis HOD model would need to improve in order to obtain a better description of the data, especially for clustering. We have found that fitting $\gamma_{t}$ only results in much better $\chi^{2}$ values than jointly fitting with $w$. However, clustering controls the satellite fraction better, especially for high-mass galaxies. In more detail, at high-mass, high-redshift bins where the contribution from the 1-halo central term is smaller, the 2-halo central and satellite terms can become degenerate, and thus the fit cannot distinguish between the two components. We have seen something similar in Zacharegkas et al. (2022) where we imposed a prior on the satellite fraction. In this work, clustering helps to control the fraction of satellites due to how sensitive $w$ is to the satellite terms.
@@
In addition, we notice that bins =$\ell=\{1,2\}$ look qualitatively different from the last two, the former having a narrower central occupation component, and a steeper satellite term. The narrower central HOD in the lower-redshift bins can be understood by the fact that at lower redshift the presence of the 1-halo term is stronger and we can thus constrain the HOD a lot better. This is both due to the fact that as the distance to the lenses increases with redshift, a given physical scale corresponds to a larger scale, as well as due to the scale cuts that remove more of the small scales at larger redshift. The steeper satellite fraction at higher halo masses corresponds to the fact that higher-mass halos host more satellite galaxies compared to the less massive halos.
@@
Fiducial $11.559^{+0.334}_{-0.415}$ $-1.689^{+0.333}_{-0.220}$ $-1.637^{+0.107}_{-0.096}$ $0.588^{+0.265}_{-0.220}$ $4.227^{+2.223}_{-1.776}$
Including highest-mass bins $11.485^{+0.287}_{-0.366}$ $-1.722^{+0.298}_{-0.211}$ $-1.645^{+0.106}_{-0.099}$ $0.571^{+0.222}_{-0.220}$ $4.326^{+1.703}_{-1.512}$

### v3

@@ 4.1.2 The stellar mass sample
With the GalPro output from our approximately complete galaxy sample, we now define the stellar mass sample for this work. In particular, we like to define bins in both redshift and stellar mass, as we are interested in the redshift-dependence of the SHMR. For the redshift binning, we define the binning similar to that used in DES Y3 (e.g. Prat2022; y3-3x2pt). That is, we bin the galaxies into four tomographic redshift bins based on their mean redshift $z_{\rm m}$, as shown in Table 1. For each redshift bin, we first impose a stellar mass cut on the mean stellar mass $M_{\star,\rm m}$ so that the remaining galaxies are complete in stellar mass at that redshift [^1] This is done using a similar procedure as that is described in Section 5.2 in Pozzetti2010. For the four redshift bins we get $\log M_{\rm lim}=(9.5617,9.9996,10.3549,10.5923)$.. We then split the remaining galaxies into stellar mass bins again according to their mean stellar mass $M_{\star,\rm m}$ where aside from the highest mass bins, each bin contains $3\times 10^{6}$ objects. The number of objects, number density, as well as the limits in stellar-mass per bin are shown in Table 1. From now on, we will denote a single bin with the indices $(\ell,m)$, the first of which refers to the reshift bin and the latter to the stellar-mass bins in that $\ell$. The approximate redshift bin edges are defined, per $\ell$ index, as $\{(0.20,0.40),(0.40,0.55),(0.55,0.70),(0.70,0.85)\}$ for $\ell=\{1,2,3,4\}$ respectively.
@@
($\log M_{\star}/M_{\odot}$) $\times 10^{6}] [$$\rm arcmin^{-2}$}] $M_{\odot}]$
@@ 4.1.3 Weights
Thus far, we have constructed a sample of galaxies that is complete in stellar mass and binned by stellar mass and redshift. However, in most cases, the rather simplistic selection will still leave low-level spatial patterns in the galaxy distributions that are imprints of observing conditions such as depth or seeing. A common way to further remove these effects is to assign each galaxy a weight to compensate the over- or under-estimation of the galaxy densities – we will refer these weights as large-scale structure (LSS) weights (see also Rodriguez-Monroy2022; HighzSample).
@@
The methodology we implement in this work is very similar to what is done in HighzSample. It relies on using a neural network to find the mapping between the galaxy number density and a number of survey conditions, and then correcting for them, with the underlying assumption that the cosmological LSS signal should not be correlated with observing conditions. In particular, we examine the following observational parameters via a set of survey property (SP) maps, each constructed in all four grizgriz bands:
@@
  • Galactic extinction: We use the SFD dust extinction map from schlegel98, which measures the E(BV)E(B-V) reddening, in magnitudes.
  • Stellar density: We use a map of stellar density, in ${\rm deg}^{-2}$, using stellar sources from Gaia EDR3 (gaia21).
@@ 4.2 Source galaxies
We use the DES Y3 shear catalog presented in y3-shapecatalog. The galaxy shapes are estimated using the Metacalibration (huff2017; sheldon2017) algorithm. The shear catalog has been thoroughly tested in y3-shapecatalog, and tests specifically tailored for tangential shear have been presented in Prat2022. In this paper, we perform additional tests on this shear catalog for tangential shear measurement on small scales (Section 5.3).
We further note that, it was recently discovered that the Year 3 DES cosmological inference results in DESY3_cosmological_inference used a tomographic binning of the source catalog that was inconsistent with the redshift distribution of the galaxies. The first iteration of our results also suffered from this inconsistency but we have since updated all measurements and our Bayesian analysis to use the correct version of the source catalog, also used in McCullough2024. Comparing the two analyses, we found mostly negligible effects from updating the catalog, although some of the trends in our host halo mass, galaxy bias and satellite fraction constraints changed as a function of stellar mass and redshift. We only present results using the correct version of the catalog in this current paper.
Following Prat2022 we bin the source galaxies into four redshift bins, where details of the redshift binning and calibration is described in y3-sompz. The redshift distributions of the sources is shown in Figure 3. The number of galaxies per source bin and the galaxy number density per square arcmin are both listed in Table 1. The approximate redshift bin edges, per $s$ index, defined as $\{(0.20,0.43),(0.43,0.63),(0.63,0.90),(0.90,1.30)\}$ for $s=\{1,2,3,4\}$ respectively.
@@ 5.1 Data vectors
The galaxy-galaxy lensing estimator is the same as in Prat2022 and written as
@@
For galaxy clustering we adopt the Landy-Szalay (landy93) estimator, written as:
@@
For both $\gamma_{t}(\theta)$ and $w(\theta)$ we use $N_{\theta}=30$ angular bins in the range of $0.25,250]\;{\rm arcmin}$. Each galaxy $i$ in our sample is weighted by $w_{g,i}$ as described in Section 4.1.3. Our catalog of random points contains $30$ times the number of galaxies in our lens galaxy catalog.
@@ 5.2 Covariance matrix
where $\gamma_{t}^{k}(\theta_{i})$ is the shear in the $i$’th angular bin for the $k$’th JK resampling, $\langle\gamma_{t}(\theta_{i})\rangle_{k}$ is the average over all $N_{\rm JK}$ realizations of the shear for the $i$’th angular bin and we have defined $\Delta\gamma^{k}_{i}\equiv\gamma_{t}^{k}(\theta_{i})-\langle\gamma_{t}(\theta_{i})\rangle_{k}$. We use $N_{\rm JK}=200$ JK patches defined via the kmeans [^2] https://github.com/esheldon/kmeans_\_radec algorithm. $N_{\rm JK}$ is chosen so that the individual JK regions are at least as large as the maximum angular scale we need for our measurements.
When inverting the covariance matrix in the likelihood analysis, a correction factor is needed to account for the bias introduced from the noisy covariance (Friedrich2016). This correction is often referred to as the Hartlap (hartlap2007) correction. When inverting the JK covariance matrix $\mathcal{C}$ we multiply it by a factor $H$ to get the unbiased covariance (kaufmann1967)
@@
where $N_{\theta}$ is the number of angular bins we use in each model fit. As shown in hartlap2007, for $N_{\theta}/N_{\rm JK}<0.8$ the correction produces an unbiased estimate of the inverse covariance matrix; in our case we find that, depending on the bin we consider, $N_{\theta}/N_{\rm JK}$ can vary between $0.215$ and $0.240$. In Appendix D we also compare our Jackknife covariance with an analytical Gaussian covariance calculated via the software package CosmoCov (Krause2017_cosmolike; fang20_cosmocov2) to show that they are fairly consistent and that our main results should not be sensitive to our choice of how we model the covariance matrix.
@@ 5.4 Bin selection and scale cuts
$\Delta z_{s}^{3}$ $\mathcal{N}(0,0.01)$ Source photo-$z$ uncertainty in 33rd bin; eq. (32)
@@
$m_{s}^{3}$ $\mathcal{N}(-0.0255,0.0013)$ Multiplicative bias in 33rd source bin; eq. (27)
@@ 6 Parameter inference
We continue to perform parameter inference by fitting our model described in Sections 2 and 3 to the data vector described in Section 5. In our fiducial model, we vary the seven SHMR parameter vector $\boldsymbol{\theta}_{\rm SHMR}=(\log M_{1},\log\epsilon,\alpha,\gamma,\delta,B_{\rm sat},B_{\rm cut},\beta_{\rm sat},\beta_{\rm cut})$. We also vary three HOD parameters $\boldsymbol{\theta}_{\rm HOD}=(\sigma_{\log M_{\star}},\alpha_{\rm sat},\kappa)$, as well as the additional model-complexity parameters $\boldsymbol{\theta}_{\rm complex}=(\kappa_{c},\alpha_{\rm lmag})$. Since the distribution of satellite galaxies is typically less concentrated than that of the dark matter (e.g. Carlberg1997; Lin2004; Nagai2004; Hansen2004), we vary $\kappa_{c}$ within a range mostly below unity. For the magnification parameter $\alpha_{\rm lamg}$ we use values between $1$ (for no magnification) and $3$ in order to be broadly consistent with the constraints for the MagLim sample from DES Y3 (DESY3_magnification). Given that our samples are expected to be complete in our redshift-stellar mass bins, we decided to set the central and satellite parameters $f_{\rm cen}$ and $f_{\rm sat}$, respectively, to unity.
@@
For observational systematic effects, we further vary a set of systematic parameters for each lens-source combination $\boldsymbol{\theta}_{\rm sys}=(m_{s}^{j},\Sigma_{\ell}^{i},\Delta z_{\ell}^{i},\Delta z_{s}^{j},\Delta\log M_{\star}^{i})$ for lens bin $i$ and source bin $j$. The functions for all these uncertainty parameters are defined in Section 3.5. To the source-bin systematics parameters $\Delta z_{s}$ and $m_{s}$ we apply the same Gaussian priors as in the DES Y3 main analyses (see Myles2021; MacCrann2022). The priors on the lens-bin redshift uncertainty parameters $\Delta z_{\ell}$ and $\Sigma_{\ell}$ are set so that they are broadly consistent with the MagLim sample from DES Y3. More specifically, we apply a Gaussian prior on $\Delta z_{\ell}$ centered at 0 with a variance of $0.1$ and a uniform prior from $0.8$ to $1.3$ to the stretch parameter $\Sigma_{\ell}$. Finally, we allow the stellar-mass distribution of our lenses to have a small shift, $\Delta\log M_{\star}$, to which we apply uniform priors from 0.15-0.15 to $0.15$, to account for about $0.3$ dex uncertainty on the mean stellar mass which is typical for inference of this parameter from SED fitting algorithms. This is a fairly conservative choice given our validation tests on the stellar mass catalog (Appendix B).
@@
The priors on the lens magnification parameter $\alpha_{\rm lmag}$ is allowed to vary from $1$, which translates to no magnification, to $3$, which gives a lot of freedom to this parameter component. The parameter $\kappa_{c}$, described in Section 3.6, is allowed to mostly get values less than unity since the satellite galaxies are expected to be less concentrated than central. The SHMR parameter priors are informed by Kravtsov2018 and allow for considerable freedom during our model fits. As for the HOD parameters, we set broad priors similar to relevant studies, such as leauthaud2011; zu2015.
@@
To sample the posterior of each data set we utilize the Multinest [^3] https://github.com/JohannesBuchner/MultiNest sampler, which implements a nested sampling algorithm (see for example feroz2009), and assume a Gaussian likelihood. For data vector $\mathbf{d}$ of length $N_{d}$ and model prediction vector $\mathbf{m}$ of the same length we express the log-likelihood as
@@
where $\mathcal{C}_{\Pi}$ and $\mathcal{C}_{p}$ are the prior and posterior covariance, respectively. We should note here that a goodness-of-fit estimation based on finding an effective number of parameters is not always straightforward when the parameters do not enter the model linearly, as discussed in Section 6.3 of Joachimi2021.
@@ 7 Results
We now present the main results from this work – the model fits to our data as well as constraints on the physical quantities of our model (galaxy bias, satellite fraction, and the SHMR). We first discuss our fiducial results in Section 7.1 and then explore some of the dependencies of our model fit to some analysis choices in Section 7.2. We use our pipeline to jointly fit lensing and galaxy clustering. As we have previously shown in Zacharegkas2022, our HOD model is capable of describing galaxy-galaxy lensing well. The two main advances compared to that work are: 1) we are using the stellar mass-selected generic galaxy sample described in Section 5, and 2) we include galaxy clustering as part of the model fit.
@@ 7.1 Constraints from DES Y3 galaxy clustering and galaxy-galaxy lensing
For all the fits, we show the model broken down into the 1- and 2-halo, central and satellite terms. We find that our model generally provides a good enough description of the data points, in all lens and source pairs and on all scales, to capture the main features of the dataset. However, our model is not a perfect fit to the data, especially for $w(\theta)$. In particular, we find that to fit the 1-halo to 2-halo transition regime for low-mass low-redshift bins, it is necessary to use the prescription described in Section 3.3 instead of the model used in Zacharegkas2022.
We have used the reduced $\chi^{2}$ values, which we list in Table 1, from fitting our model to the data as a goodness-of-fit guide when determining how well the fits perform using different iterations of our fiducial model. The goodness-of-fit numbers we obtain are much larger than unity per degree of freedom, which, at first glance, implies a poor fit of the model to the data. In principle, we could add additional complexities (see Section 3.6) to the model to make the $\chi^{2}/N_{\rm dof}$ closer to 1, but that would create additional degeneracies that can be difficult to interpret. Rather, we keep the most relevant and physically motivated model components here to capture the qualitative behaviors of the data. In future work, the basis HOD model would need to improve in order to obtain a better description of the data, especially for clustering. We have found that fitting $\gamma_{t}$ only results in much better $\chi^{2}$ values than jointly fitting with $w$. However, clustering controls the satellite fraction better, especially for high-mass galaxies. In more detail, at high-mass, high-redshift bins where the contribution from the 1-halo central term is smaller, the 2-halo central and satellite terms can become degenerate, and thus the fit cannot distinguish between the two components. We have seen something similar in Zacharegkas2022 where we imposed a prior on the satellite fraction. In this work, clustering helps to control the fraction of satellites due to how sensitive $w$ is to the satellite terms.
@@
In addition, we notice that bins $\ell=\{1,2\}$ look qualitatively different from the last two, the former having a narrower central occupation component, and a steeper satellite term. The narrower central HOD in the lower-redshift bins can be understood by the fact that at lower redshift the presence of the 1-halo term is stronger and we can thus constrain the HOD a lot better. This is both due to the fact that as the distance to the lenses increases with redshift, a given physical scale corresponds to a larger scale, as well as due to the scale cuts that remove more of the small scales at larger redshift. The steeper satellite fraction at higher halo masses corresponds to the fact that higher-mass halos host more satellite galaxies compared to the less massive halos.
@@
Fiducial $11.506^{+0.325}_{-0.404}$ $1.632^{+0.306}_{-0.181}$ $1.638^{+0.108}_{-0.099}$ $0.596^{+0.251}_{-0.210}$ $3.810^{+2.045}_{-1.811}$
Including highest-mass bins $11.485^{+0.287}_{-0.366}$ $1.722^{+0.298}_{-0.211}$ $1.645^{+0.106}_{-0.099}$ $0.571^{+0.222}_{-0.220}$ $4.326^{+1.703}_{-1.512}$